LARGE AMPLITUDE MOTIONS IN COUMARAN: A SPECTROSCOPIC AND THEORETICAL STUDY

\underline{PAOLO OTTAVIANI}, W CAMINATI, Dipartimento di Chimica “G. Ciamician” Università di Bologna, Via Selmi 2, I-40126 Bologna, Italy.

The structure of coumaran (2,3-dihydrobenzofuran) is characterized by a non-planar skeleton, due to the position of the most far apart carbon in the dihydrofuran ring. We have measured 133 a, b and c type lines of the rotational spectrum in the millimeter-wave region (59-78GHz), characterized by a doubling typical of an inversion motion such as the ring-puckering of the five-membered ring. By using a coupled Watson’s Hamiltonian we found an energy splitting E= 93682.019(22) MHz; this value is quite in contrast with a previous far-infrared and Raman study^{a} . In order to better interpret this large amplitude motion we performed also quantum chemical calculation (MP2 6-311++G** and B3LYP 6-311++G**) and used Meyer’s flexible model.

aE. Bondoc, T. Klots, J. Laane, J. Phys. Chem A 104, 275 (2000)